Q:The fourth term of the geometric progression is 8 and the sum of first 6 terms of the series is 63. Find the first term and the common difference?
Answer by richwmiller(17219)

Tn = t * r^(n - 1)
8= t * r^3

S=t*(1 - r^n)/(1 - r)
63=t*(1 - r^6)/(1 - r),
8= t * r^3
common difference r=2, first term t=1.
1,2,4,8,16,32
1+2+4+8+16+32=63